OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (nirmalnema):

|x − 2| + 10 = 12 |x − 2| = 2 if x<0 |x − 2| is -ve -x+2=2x=0 |x − 2| is +ve x-2=2 x=4 x=0,4